OpenStudy (anonymous):

10x2 + 7x + 2, when x = -2

OpenStudy (anonymous):

\[10x^2 + 7x + 2\]\[10\times (-2)^2+7\times (-2)+2\]

OpenStudy (anonymous):

then 10*4-14+2=28